A man is nine times as old as his son. In nine years, he will be only three times as old as his son. How old are each now?
Solve using elimination method.

Answer by mananth(16949)

let son be x years
father = y
...
y=9x
after 3 years
son = x+9
father = y+9
(y+9)=3(x+9)
plug value of y =9x
(9x+9)=3x+27
9x+9=3x+27
-3x
9x-3x+9=27
6x+9=27
-9
6x=18
/6
x=3 son's age now
father's age = 9x = 27 years now
..
m.ananth@hotmail.ca